OpenStudy (anonymous):

I need help to; Use intergration by parts to evaulate the integral ∫ (x) √4:(x+7)dx

OpenStudy (anonymous):

\[x \sqrt[4]{x+7}\]

OpenStudy (zarkon):

u=x+7

OpenStudy (zarkon):

no need to use parts

OpenStudy (anonymous):

Why wouldn't you use parts

OpenStudy (zarkon):

if u=x+7 then x=u-7 du=dx \[\int x \sqrt[4]{x+7}dx=\int(u-7)u^{1/4}du=\int\left(u^{5/4}-7u^{1/4}\right)du=\cdots\]